# Tunneling in double-well potentials within stochastic quantization: Application to ammonia inversion

### v1

@@ III Double Square Well
where L=(bd)/2L=(b-d)/2. The transcendental equation (46) thus determines the allowed energies (i.e., the wavevectors $k$) for the symmetric states.
@@
Similarly, the solutions for the antisymmetric (odd) wavefunctions are of the form
@@ III.1 Tunneling Time for the Ground State
In Fig. 2, we present, in a semi-logarithmic plot, the behavior of $\bar{\tau}$ given by Eq. (50) as a function of $k_{0}$ (recall that $k_{0}=\sqrt{2mV_{0}}/\hbar$) for three values of $d$. Here we plot $\bar{\tau}$ in dimensionless units, with $\hbar=m=1$ and $b=6$, for $d=2.0$, $3.0$, and $4.0$. For each set of parameters, we solve the transcendental equation (46) to determine the lowest $k$, so we can compute $\bar{\tau}$ from Eq. (50). We see from this figure that $\bar{\tau}$ increases monotonically with $k_{0}$, as expected, and exhibits an exponential growth at large $k_{0}$, in agreement with Eq. (51)—a behavior clearly reflected by the linear asymptotic trend in the semi-logarithmic plot. Figure 3 shows $\bar{\tau}$ as a function of $d$, for three different values of $V_{0}$. One sees from this figure that, starting from a small value of $d$, the mean tunneling time first increases with the barrier width, as expected, then reaches a maximum value, and decreases after that. This behavior is explained by the fact that increasing $d$, for fixed $b$, leads to a reduction in the distance, $L$, between the impenetrable walls and the barrier, see Fig. 1, so that for sufficiently small $L$ the particle becomes ‘compressed’ in a narrow region between the infinite wall and the barrier wall. In such a situation, the particle ‘ricochets’ ever so more rapidly between these two vertical walls, which leads to an increasing (decreasing) of the tunneling probability (time).
@@
We briefly comment on the tunneling process from an energetic viewpoint. We recall that in SQ the instantaneous energy of the particle is given by [Nelson1979]
@@
where $\rho(x)=|\psi_{0}(x)|^{2}$ and $\psi_{0}(x)$ is the ground-state wavefunction. Equation (53) shows that, within the barrier region, the instantaneous stochastic energy always satisfies ${\cal E}(x)\geq V_{0}$. Thus, strictly speaking, the standard textbook picture of quantum tunneling—where the particle traverses a classically forbidden region with insufficient energy—does not apply in the SQ framework. Instead, the stochastic dynamics continually exchanges energy with the background noise, enabling the particle to transiently acquire enough energy to cross the barrier, even though the mean value $E_{0}$ remains below the barrier height.
Figure 5: Sample time series of the particle’s instantaneous energy ${\cal E}(x)$ in the double square-well potential obtained via stochastic quantization. The mean energy $\expectationvalue{{\cal E}(x)}$ coincides with the ground-state eigenenergy $E_{0}$ of the system. The inset displays the corresponding stochastic trajectory, with the barrier edges indicated by the dotted lines. The instantaneous energy exceeds the barrier height $V_{0}$ whenever the particle is inside the barrier region, see, e.g., the highlighted region. The instantaneous energy also exhibits a sharp rise whenever the particle tries to approach the infinite walls of the potential.
@@ III.2 Quantum Mechanics Prediction: Probability Oscillation Period
The situation is fundamentally different in Nelson’s stochastic quantization, where the probability density arises from an underlying stochastic process. In this framework, a mean tunneling time, $\bar{\tau}$, can be computed even for stationary states. Despite this conceptual difference, we shall show below that there exists a direct quantitative relation between the mean tunneling time predicted by the stochastic quantization formalism and that inferred from the two-state quantum-mechanical description.
@@ IV Tunneling Time for Generic Double-Well Potentials: WKB Analysis
For a continuous potential, it is not straightforward to identify where the particle begins and ends the tunneling process, since the barrier does not necessarily have sharp boundaries. Köppe et al[KoeppeJMP2018_QHE_1D] suggested that one can consider the traversal time that a particle, initially released at some point $x$ in the left well, i.e., $\infty<x<-b$, would take to reach the right classical turning point $x=b$, and then average over the initial positions. However, for a high barrier, i.e., $V_{0}\gg E_{0}$, the barrier walls become sufficiently sharp that one can assume, as was the case of the double square well potential, that tunneling effectively starts when the particle reaches the right classical turning point, x=bx=-b, of the left well and ends when the particle ‘emerges’ at the symmetric point, $x=b$, in the right well; see Fig. 6. Also, because the potential is assumed to have unbounded support, we take $a=$-$\infty$, corresponding to a reflecting boundary at minus infinity; see Sec. II.2. With these choices, then according to Eq. (38) the mean tunneling time is given by
@@
We recall that the WKB approximation [LandauLifshitzQM, MerzbacherQM] is valid when the potential changes very slowly compared to the particle wavelength, which implies the condition:
@@
Using the WKB connection formulas for matching wave functions across turning points [LandauLifshitzQM, MerzbacherQM], the symmetric wave function can be written in the relevant regions as
@@
Let us now turn to the probability oscillation period in the framework of quantum mechanics. It is well known [LandauLifshitzQM, David-Park] that in the WKB approximation the energy splitting for a symmetric double-well in the high barrier limit ($\Phi\gg 1$) is given by
@@
It is thus noteworthy that both the mean tunneling time and the quantum-probability oscillation period have the same dependence on the barrier integral $\Phi$ and on the classical time $T_{\text{cl}}$, differing only by a numerical factor; see Eqs. (80) and (82). From these relations we then immediately recover relation (63) for the ratio (in the high-barrier limit) between the tunneling time, $\tau_{\rm QM}=T/2$, estimated from quantum mechanics and the mean tunneling time, $\bar{\tau}$, predicted by stochastic quantization. Note, however, that while Eq. (63) was derived for the double square well, the WKB calculation presented above is valid regardless of the specific shape of the potential or energy level in the WKB limit. Both calculations yield the same numerical factor between $\tau_{\rm QM}$ and $\bar{\tau}$; yet it remains unclear whether the value $\pi/2$ admits a simple physical interpretation.
@@ V Application to Ammonia
A well-known example of quantum tunneling is the inversion of the ammonia molecule (NH3). In this process, the nitrogen atom can penetrate the potential barrier formed by the plane of the three hydrogen atoms, effectively oscillating between two equivalent configurations on either side of the plane. Within quantum mechanics, this inversion is understood as the time evolution of a localized state that arises from the coherent superposition of the system’s symmetric (even) and antisymmetric (odd) energy eigenstates. The transition frequency—here denoted by $\nu_{\mathrm{QM}}$—for such localized states is related to the energy splitting $\Delta E$ between the ground and first excited states of ammonia, or more exactly $\nu_{\rm QM}={\Delta E}/{2\pi\hbar}$. The experimental value for the ammonia inversion frequency is approximately $24\,\text{GHz}$ [TownesSchawlow1955].
In this section, we shall compute the mean tunneling time $\bar{\tau}$ predicted by Nelson’s stochastic quantization for the ground state of the ammonia molecule by using suitable double-well potentials. Using the correspondence given in Eq. (63) between the mean tunneling time, $\bar{\tau}$, obtained from stochastic quantization and the oscillation time, $T$, computed from quantum mechanics, we can then write a stochastic-quantization prediction for the ammonia inversion frequency as follows:
$\nu_{\rm SQ}\equiv\frac{1}{\pi\bar{\tau}}.$ (83)
As argued above, $\nu_{\rm SQ}$ should in general be relatively close to the inversion frequency, $\nu_{\rm QM}$, predicted by quantum mechanics. Indeed, we will see below that the prediction for the inversion frequency defined in Eq. (83) agrees remarkably well with the experimental value for the ammonia inversion frequency, thus validating the calculation of the mean tunneling time within the stochastic quantization formalism.
To investigate the tunneling dynamics of the ammonia molecule within the framework of stochastic quantization, we model the system as a particle of mass mm—the reduced mass of the molecule—moving in a suitable symmetric double-well potential. Several potential models have been proposed to describe the inversion motion in ammonia, including the Dennison–Uhlenbeck potential [DennisonUhlenbeck1932], the Rosen–Morse potential [RosenMorse], and the Manning potential [Manning], among others [Peacock-Lopez, Sitnitsky2017]. However, as discussed above, the near identity between the tunneling frequency predicted by stochastic quantization, $\nu_{\rm SQ}$, and its quantum-mechanical counterpart, $\nu_{\rm QM}$, is expected to hold irrespective of the detailed form of the potential (in the high-barrier limit). Hence, for our purposes it suffices to consider a representative potential model for the ammonia molecule. In what follows, we therefore adopt the double Rosen–Morse potential (RMP), which has been widely used to model the ammonia inversion dynamics [RosenMorse, YangHan_IJMS_2021_Ammonia].
@@
The double Rosen-Morse potential was introduced by Rosen and Morse [RosenMorse] to model the vibration of polyatomic molecules, with special attention to the ammonia inversion problem, which is the prototypical example of this kind of problem. It has the following form (see Fig. 7):
@@
To choose the potential parameters one typically uses three experimental data [DennisonUhlenbeck1932], namely: the energy separation, $E_{1,+}-E_{0,-}=950\text{ cm}^{-1}$, between the two lowest pair of levels and the separations within each pair of levels, $\Delta E_{0}=E_{0,+}-E_{0,-}=0.8\text{ cm}^{-1}$ and $\Delta E_{1}=E_{1,+}-E_{1,-}=33\text{ cm}^{-1}$ (where $1\text{ cm}^{-1}=1\,hc\text{ J}\approx 1.988\cdot 10^{-23}\text{ J}$), see Fig. 7. Since there are four parameters and only three experimental data to fit, the best that can be done in this case is to determine a range of values for each parameter based on the reasonable shape of the potential, as described in [RosenMorse]. These restrictions provide the following ranges of possible values for the RMP parameters: $0<A<1000\text{ cm}^{-1},\ 2200\text{ cm}^{-1}<B<3000\text{ cm}^{-1},\ 0.16\text{ \AA }<d<0.185\text{ \AA }$, and $2.20<k<2.24$.
@@
Within the ranges of allowed parameters mentioned above, one can verify that the values of $x_{0}$ and $V_{0}$ differ at most by eight percent. So, any form of the potential (84) within the acceptable range of parameters will have $x_{0}$ near 0.38 Å and $V_{0}$ near 2050 cm-1. With these considerations in mind, we choose the values of the parameters in order to get a value of $\Delta E_{0}$ very close to the experimental value of $0.8\text{ cm}^{-1}$, since this is the energy splitting that determines the quantum-mechanical tunneling frequency. To do so, we compute numerically the energy eigenvalues of the RMP with the Numerov method [Numerov1924] and apply a fit procedure to determine the optimal RMP parameters, where we fix $d=0.17\text{ \AA }$ and $k=2.22$, and vary $A$ and $B$. The values obtained are as follows: $A=398\text{ cm}^{-1}$ and $B=2810\text{ cm}^{-1}$, which of course are within the allowed ranges, as expected.
Having determined the relevant parameter values for the double Rosen-Morse potential, we can proceed to compute the mean tunneling time predicted by stochastic quantization for the ground state of the ammonia molecule. The classical turning points are $x=\pm b$ and $x=\pm c$, where $V(\pm c)=V(\pm b)=E_{0,-}$. The ground state wave function, $\psi_{0}(x)$, and its energy, $E_{0,-}$, were calculated numerically using the Numerov method. We then numerically performed the integrations in Eq. (64) to obtain the mean tunneling time, $\bar{\tau}$, for a particle with reduced mass $m$ given by Eq. (85) in the ground state of the RMP. We find $\bar{\tau}=13.4578$ ps, which in view of Eq. (83) corresponds to a tunneling frequency $\nu_{\rm SQ}=23.65$ GHz, which is indeed very close to the experimental value of 23.79 GHz [TownesSchawlow1955]. This agreement is quite remarkable, when considering that the ‘conversion formula’ (63) between tunneling time and quantum oscillation period, which led to Eq. (83), was derived in the asymptotic limit $V_{0}\gg E$. Our results thus suggest that this correspondence is expected to hold for any practical tunneling situation where the quantum mechanical two-state approximation is valid, as in the inversion dynamics of the ammonia molecule.
@@
In Fig. 8 we show the distribution of tunneling times for the ground state of the RMP with the parameters specified above, obtained from simulations of $10^{6}$ trajectories of the Brownian particle evolving under the corresponding stochastic quantization diffusion process. The simulations were carried out using the Euler–Maruyama method with a time step of $10^{-3}$ in dimensionless units (or approximately $10^{-5}\,\text{ps}$). For each trajectory, the particle was initialized at $t=0$ at the right classical turning point of the left potential well, and the simulation was terminated when the particle reached the left classical turning point of the right well. The elapsed time between these two events defines the tunneling time for the trajectory. As seen in Fig. 8, the resulting distribution again exhibits a clear exponential decay, in agreement with the theoretical prediction discussed in Sec. II.2. Note also that, in contrast to the corresponding distribution for the double square well (see Fig. 4), the RMP case exhibits a much sharper rise from $p(0)=0$ to the exponential curve for $\tau>0$, with no events outside the exponential region detected within our set of $10^{6}$ simulations. Consequently, an exponential fit (red line) provides an excellent description of the entire dataset. The theoretical mean tunneling time, computed from Eq. (64), is $\bar{\tau}_{\rm th}=13.46~\text{ps}$, in very good agreement with the corresponding numerical value, $\bar{\tau}_{\rm num}=13.86~\text{ps}$.

### v2

@@ III Double Square Well
where L=(bd)/2L=(b-d)/2. The transcendental equation (46) thus determines the allowed energies (i.e., the wavevectors $k$) for the symmetric states. We note that for fixed $L$, the minimum barrier height for which a bound state still exists with $E<V_{0}$ is
$V_{0,\min}=\frac{\hbar^{2}\pi^{2}}{8mL^{2}}\,$, (47)
corresponding to the situation when the bound state energy $E$ just touches the top of the barrier. Similarly, for fixed $V_{0}$, the maximum barrier width that allows for a bound state with $E<V_{0}$ is
$d_{\max}=b-\frac{\pi\hbar}{\sqrt{2mV_{0}}}\,.$ (48)
@@
The solutions for the antisymmetric (odd) wavefunctions are of the form
@@ III.1 Tunneling Time for the Ground State
In Fig. 2, we present, in a semi-logarithmic plot, the behavior of $\bar{\tau}$ given by Eq. (52) as a function of $k_{0}$ (recall that $k_{0}=\sqrt{2mV_{0}}/\hbar$) for three values of $d$. Here we plot $\bar{\tau}$ in dimensionless units, with $\hbar=m=1$ and $b=6$, for $d=2.0$, $3.0$, and $4.0$. For each set of parameters, we solve the transcendental equation (46) to determine the lowest $k$, so we can compute $\bar{\tau}$ from Eq. (52). We see from this figure that $\bar{\tau}$ increases monotonically with $k_{0}$, as expected, and exhibits an exponential growth at large $k_{0}$, in agreement with Eq. (53)—a behavior clearly reflected by the linear asymptotic trend in the semi-logarithmic plot. Figure 3 shows $\bar{\tau}$ as a function of $d$, for three different values of $V_{0}$. One sees from this figure that, starting from a small value of $d$, the mean tunneling time first increases with the barrier width, as expected, then reaches a maximum value, and decreases after that. This behavior reflects a competition between two effects. On the one hand, increasing $d$ widens the barrier, which tends to suppress tunneling. On the other hand, reducing the classically allowed region (i.e., decreasing $L$) increases the kinetic energy inside the wells, causing the particle to traverse the allowed region more rapidly and encounter the barrier more frequently. In other words, the particle ‘ricochets’ ever so more rapidly between these two vertical walls. Beyond a certain value of $d$ (at which $\bar{\tau}$ reaches a maximum), this confinement effect dominates, leading to a decrease in the tunneling time. This interplay explains the overall shape of the curves shown in Fig. 3. The fact that the function $\bar{\tau}(d)$, for fixed $V_{0}$, exhibits a maximum explains the curve crossings observed in Fig. 2. For example, in the case $V_{0}=2$, the maximum occurs near $d\simeq 3$ (see the black curve in Fig. 3), which accounts for the observation that around $k_{0}=2$ in Fig. 2 the curve corresponding to $d=3$ lies above those for $d=2$ and $d=4$. For sufficiently large $k_{0}$, however, the maxima of $\bar{\tau}(d)$ shift to values of $d$ beyond the three considered in Fig. 2, and the curves become monotonically ordered from bottom to top as $d$ increases.
@@
We briefly comment on the tunneling process from an energetic viewpoint. We recall that in SM the instantaneous energy of the particle is given by [44]
@@
where $\rho(x)=|\psi_{0}(x)|^{2}$ and $\psi_{0}(x)$ is the ground-state wavefunction. Equation (55) shows that, within the barrier region, the instantaneous stochastic energy always satisfies ${\cal E}(x)\geq V_{0}$. Thus, strictly speaking, the standard textbook picture of quantum tunneling—where the particle traverses a classically forbidden region with insufficient energy—does not apply in the SM framework. Instead, the stochastic dynamics continually exchanges energy with the background noise, enabling the particle to transiently acquire enough energy to cross the barrier, even though the mean value $E_{0}$ remains below the barrier height.
Figure 5: Sample time series of the particle’s instantaneous energy ${\cal E}(x)$ in the double square-well potential obtained via stochastic mechanics. The mean energy $\expectationvalue{{\cal E}(x)}$ coincides with the ground-state eigenenergy $E_{0}$ of the system. The inset displays the corresponding stochastic trajectory, with the barrier edges indicated by the dotted lines. The instantaneous energy exceeds the barrier height $V_{0}$ whenever the particle is inside the barrier region, see, e.g., the highlighted region. The instantaneous energy also exhibits a sharp rise whenever the particle tries to approach the infinite walls of the potential.
@@ III.2 Quantum Mechanics Prediction: Probability Oscillation Period
The situation is fundamentally different in Nelson’s stochastic mechanics, where the probability density arises from an underlying stochastic process. In this framework, a mean tunneling time, $\bar{\tau}$, can be computed even for stationary states. Despite this conceptual difference, we shall show below that there exists a direct quantitative relation between the mean tunneling time predicted by the stochastic mechanics formalism and that inferred from the two-state quantum-mechanical description.
@@ IV Tunneling Time for Generic Double-Well Potentials: WKB Analysis
For a continuous potential, it is not straightforward to identify where the particle begins and ends the tunneling process, since the barrier does not necessarily have sharp boundaries. Alternative approaches based on the construction of time operators, such as that of Ref. [17], address this issue in a different manner by defining a traveling time between prescribed spatial regions—whether or not tunneling occurs—through the expectation value of an appropriate self-adjoint operator [13]. Within the SM approach, Köppe et al[26] suggested instead considering the traversal time for a particle initially released at a position $x$ in the left well (i.e., $\infty<x<-b$) to reach the right classical turning point at $x=b$, and then averaging over the initial positions. However, for a high barrier, i.e., $V_{0}\gg E_{0}$, the barrier walls become sufficiently sharp that one can assume, as was the case of the double square well potential, that tunneling effectively starts when the particle reaches the right classical turning point, x=bx=-b, of the left well and ends when the particle ‘emerges’ at the symmetric point, $x=b$, in the right well; see Fig. 6. Also, because the potential is assumed to have unbounded support, we take $a=$-$\infty$, corresponding to a reflecting boundary at minus infinity; see Sec. II.2. With these choices, then according to Eq. (38) the mean tunneling time is given by
@@
We recall that the WKB approximation [30, 39] is valid when the potential changes very slowly compared to the particle wavelength, which implies the condition:
@@
Using the WKB connection formulas for matching wave functions across turning points [30, 39], the symmetric wave function can be written in the relevant regions as
@@
Let us now turn to the probability oscillation period in the framework of quantum mechanics. It is well known [30, 50] that in the WKB approximation the energy splitting for a symmetric double-well in the high barrier limit ($\Phi\gg 1$) is given by
@@
It is thus noteworthy that both the mean tunneling time and the quantum-probability oscillation period have the same dependence on the barrier integral $\Phi$ and on the classical time $T_{\text{cl}}$, differing only by a numerical factor; see Eqs. (82) and (84). From these relations we then immediately recover relation (65) for the ratio (in the high-barrier limit) between the tunneling time, $\tau_{\rm QM}=T/2$, estimated from quantum mechanics and the mean tunneling time, $\bar{\tau}$, predicted by stochastic mechanics. Note, however, that while Eq. (65) was derived for the double square well, the WKB calculation presented above is valid regardless of the specific shape of the potential or energy level in the WKB limit. Both calculations yield the same numerical factor between $\tau_{\rm QM}$ and $\bar{\tau}$; yet it remains unclear whether the value $\pi/2$ admits a simple physical interpretation.
@@ V Application to Ammonia
A well-known example of quantum tunneling is the inversion of the ammonia molecule (NH3). In this process, the nitrogen atom can penetrate the potential barrier formed by the plane of the three hydrogen atoms, effectively oscillating between two equivalent configurations on either side of the plane. Within quantum mechanics, this inversion is understood as the time evolution of a localized state that arises from the coherent superposition of the system’s symmetric (even) and antisymmetric (odd) energy eigenstates. The transition frequency—here denoted by $\nu_{\mathrm{QM}}$—for such localized states is related to the energy splitting $\Delta E$ between the ground and first excited states of ammonia, or more exactly $\nu_{\rm QM}={\Delta E}/{2\pi\hbar}$. The experimental value for the ammonia inversion frequency is approximately $24\,\text{GHz}$ [59].
In this section, we shall compute the mean tunneling time $\bar{\tau}$ predicted by Nelson’s stochastic mechanics for the ground state of the ammonia molecule by using suitable double-well potentials. Using the correspondence given in Eq. (65) between the mean tunneling time, $\bar{\tau}$, obtained from stochastic mechanics and the oscillation time, $T$, computed from quantum mechanics, we can then write a stochastic-mechanics prediction for the ammonia inversion frequency as follows:
$\nu_{\rm SM}\equiv\frac{1}{\pi\bar{\tau}}.$ (85)
As argued above, $\nu_{\rm SM}$ should in general be relatively close to the inversion frequency, $\nu_{\rm QM}$, predicted by quantum mechanics. Indeed, we will see below that the prediction for the inversion frequency defined in Eq. (85) agrees remarkably well with the experimental value for the ammonia inversion frequency, thus validating the calculation of the mean tunneling time within the stochastic mechanics formalism.
To investigate the tunneling dynamics of the ammonia molecule within the framework of stochastic mechanics, we model the system as a particle of mass mm—the reduced mass of the molecule—moving in a suitable symmetric double-well potential. Several potential models have been proposed to describe the inversion motion in ammonia, including the Dennison–Uhlenbeck potential [15], the Rosen–Morse potential [55], and the Manning potential [37], among others [54, 57]. However, as discussed above, the near identity between the tunneling frequency predicted by stochastic mechanics, $\nu_{\rm SM}$, and its quantum-mechanical counterpart, $\nu_{\rm QM}$, is expected to hold irrespective of the detailed form of the potential (in the high-barrier limit). Hence, for our purposes it suffices to consider a representative potential model for the ammonia molecule. In what follows, we therefore adopt the double Rosen–Morse potential (RMP), which has been widely used to model the ammonia inversion dynamics [55, 64].
@@
The double Rosen-Morse potential was introduced by Rosen and Morse [55] to model the vibration of polyatomic molecules, with special attention to the ammonia inversion problem, which is the prototypical example of this kind of problem. It has the following form (see Fig. 7):
@@
To choose the potential parameters one typically uses three experimental data [15], namely: the energy separation, $E_{1,+}-E_{0,-}=950\text{ cm}^{-1}$, between the two lowest pair of levels and the separations within each pair of levels, $\Delta E_{0}=E_{0,+}-E_{0,-}=0.8\text{ cm}^{-1}$ and $\Delta E_{1}=E_{1,+}-E_{1,-}=33\text{ cm}^{-1}$ (where $1\text{ cm}^{-1}=1\,hc\text{ J}\approx 1.988\cdot 10^{-23}\text{ J}$), see Fig. 7. Since there are four parameters and only three experimental data to fit, the best that can be done in this case is to determine a range of values for each parameter based on the reasonable shape of the potential, as described in [55]. These restrictions provide the following ranges of possible values for the RMP parameters: $0<A<1000\text{ cm}^{-1},\ 2200\text{ cm}^{-1}<B<3000\text{ cm}^{-1},\ 0.16\text{ \AA }<d<0.185\text{ \AA }$, and $2.20<k<2.24$.
@@
Within the ranges of allowed parameters mentioned above, one can verify that the values of $x_{0}$ and $V_{0}$ differ at most by eight percent. So, any form of the potential (86) within the acceptable range of parameters will have $x_{0}$ near 0.38 Å and $V_{0}$ near 2050 cm-1. With these considerations in mind, we choose the values of the parameters in order to get a value of $\Delta E_{0}$ very close to the experimental value of $0.8\text{ cm}^{-1}$, since this is the energy splitting that determines the quantum-mechanical tunneling frequency. To do so, we compute numerically the energy eigenvalues of the RMP with the Numerov method [47] and apply a fit procedure to determine the optimal RMP parameters, where we fix $d=0.17\text{ \AA }$ and $k=2.22$, and vary $A$ and $B$. The values obtained are as follows: $A=398\text{ cm}^{-1}$ and $B=2810\text{ cm}^{-1}$, which of course are within the allowed ranges, as expected.
Having determined the relevant parameter values for the double Rosen-Morse potential, we can proceed to compute the mean tunneling time predicted by stochastic mechanics for the ground state of the ammonia molecule. The classical turning points are $x=\pm b$ and $x=\pm c$, where $V(\pm c)=V(\pm b)=E_{0,-}$. The ground state wave function, $\psi_{0}(x)$, and its energy, $E_{0,-}$, were calculated numerically using the Numerov method. We then numerically performed the integrations in Eq. (66) to obtain the mean tunneling time, $\bar{\tau}$, for a particle with reduced mass $m$ given by Eq. (87) in the ground state of the RMP. We find $\bar{\tau}=13.4578$ ps, which in view of Eq. (85) corresponds to a tunneling frequency $\nu_{\rm SM}=23.65$ GHz, which is indeed very close to the experimental value of 23.79 GHz [59]. This agreement is quite remarkable, when considering that the ‘conversion formula’ (65) between tunneling time and quantum oscillation period, which led to Eq. (85), was derived in the asymptotic limit $V_{0}\gg E$. Our results thus suggest that this correspondence is expected to hold for any practical tunneling situation where the quantum mechanical two-state approximation is valid, as in the inversion dynamics of the ammonia molecule.
@@
In Fig. 8 we show the distribution of tunneling times for the ground state of the RMP with the parameters specified above, obtained from simulations of $10^{6}$ trajectories of the Brownian particle evolving under the corresponding stochastic mechanics diffusion process. The simulations were carried out using the Euler–Maruyama method with a time step of $10^{-3}$ in dimensionless units (or approximately $10^{-5}\,\text{ps}$). For each trajectory, the particle was initialized at $t=0$ at the right classical turning point of the left potential well, and the simulation was terminated when the particle reached the left classical turning point of the right well. The elapsed time between these two events defines the tunneling time for the trajectory. As seen in Fig. 8, the resulting distribution again exhibits a clear exponential decay, in agreement with the theoretical prediction discussed in Sec. II.2. Note also that, in contrast to the corresponding distribution for the double square well (see Fig. 4), the RMP case exhibits a much sharper rise from $p(0)=0$ to the exponential curve for $\tau>0$, with no events outside the exponential region detected within our set of $10^{6}$ simulations. Consequently, an exponential fit (red line) provides an excellent description of the entire dataset. The theoretical mean tunneling time, computed from Eq. (66), is $\bar{\tau}_{\rm th}=13.46~\text{ps}$, in very good agreement with the corresponding numerical value, $\bar{\tau}_{\rm num}=13.86~\text{ps}$.